# for the JEM-EUSO collaboration

### v1

@@ Abstract
Fluorescence telescopes are important instruments widely used in modern experiments for registering ultraviolet radiation from extensive air showers (EASs) generated by cosmic rays of ultra-high energies. We present a proof-of-concept convolutional neural network aimed at reconstruction of energy and arrival directions of primary particles using model data for two telescopes developed by the international JEM-EUSO collaboration. We also demonstrate how a simple convolutional encoder-decoder can be used for EAS track recognition. The approach is generic and can be adopted for other fluorescence telescopes.
@@ 1 Introduction
Registering fluorescence radiation of extensive air showers (EASs) with dedicated telescopes in nocturnal atmosphere of Earth is an established technique of studying ultra-high energy cosmic rays ($\gtrsim 10^{18}\leavevmode\nobreak\ \text{eV}=1\leavevmode\nobreak\ \text{EeV}$[1]. Fluorescence telescopes (FTs) play an important role in both leading experiments in this field of astrophysics, the Pierre Auger Observatory [2] and the Telescope Array [3]. They were also suggested to be used as orbital telescopes due to the opportunity to increase the exposure of an experiment dramatically [4]. The international JEM-EUSO collaboration is implementing a step-by-step program aimed at performing a full-scale orbital experiment aimed at solving the long-standing puzzle of the nature and origin of the highest energy cosmic rays [5, 6, 7]. Two of the instruments built on this path are a small ground-based EUSO-TA telescope [8, 9, 10], which operates at the site of the Telescope Array experiment, and a stratospheric experiment EUSO-SPB2, which included both fluorescence and Cherenkov telescopes [11, 12].
Reconstruction of parameters of initial UHECRs basing on the signal registered with a monocular FT is not a trivial task [13, 14, 15]. Here we suggest two simple artificial neural networks that can be used to recognize tracks registered by a single FT and to reconstruct energy and arrival directions (ADs) of ultra-high energy primaries. We employ data simulated for EUSO-TA and the fluorescence telescope of EUSO-SPB2 to train and test these networks. In what follows, we present preliminary results of this technique and outline possible ways of its improvement. We believe that the suggested approach is generic and can be adopted for other fluorescence telescopes operating nowadays as well as the future ones. It can also be modified to use benefits provided by stereoscopic observations.
@@ 2 Fluorescence telescopes EUSO-SPB2 and EUSO-TA
EUSO-SPB2 (the Extreme Universe Space Observatory on a Super Pressure Balloon 2) was a stratospheric pathfinder mission for a future orbital experiment like POEMMA [16]. The scientific equipment included two telescopes with an identical modified Schmidt design of optics with a 1 m diameter entrance pupil occupied by an aspheric corrector plate, a curved focal surface (FS), and a spherical primary mirror. The Cherenkov telescope was aimed for observing Cherenkov emission of cosmic ray EAS with energies above 1 PeV. It could be tilted up and down between $3.5^{\circ}$ and $-13^{\circ}$ w.r.t. the horizon pointing towards the limb [17, 18]. The fluorescence telescope was aimed to register the fluorescence emission of extensive air showers born by UHECRs with an energy above 2 EeV. The telescope was pointed in the nadir direction and had a total field of view (FoV) of $12^{\circ}\times 36^{\circ}$. The FT camera consisted of three photo detector modules (PDMs), each built of $48\times 48$ pixels in a rectangular grid, thus giving 6912 pixels in total. Every PDM was covered with a BG3 filter to limit its sensitivity to the near-UV range (290–430 nm). The time resolution was equal to 1 $\mu$[19, 20].
@@
EUSO-TA can operate at different elevation angles but we are using only that of $10^{\circ}$ below. Studies for different elevation angles are to be performed in the future.
@@ 3 Simulated data sets
We used CONEX [21] to simulate EAS from UHECRs in the energy range from a few EeV up to 100 EeV. The response of both instruments was simulated with the $\overline{\rm Off}$$\underline{\rm line}$ framework [22]. For EUSO-TA, we demanded that EAS cores were within the projection of the telescope FoV on the ground. The distance between the telescope and shower cores varied in the range 2–40 km depending on the energy of primary protons. For both instruments, we used events that flagged software triggers of the respective telescope. Out of these events, we only excluded those that contained just a few hit pixels in one of the corners of the field of view of a telescope thus not providing enough data for energy or AD reconstruction. No other quality cuts on the signals were implied. In particular, we did not select events with the shower maximum being in the FoV of a telescope, which can considerably improve the quality of reconstruction.
@@ 4.1 EUSO-SPB2
Figure 1: Examples of integral tracks of simulated EASs produced by UHECRs as seen by EUSO-SPB2. A track shown in the left panel was produced by a shower initiated by a 96 EeV proton arriving at the zenith angle $\theta=19^{\circ}$. The right panel shows a track from a 26 EeV proton arriving at $\theta=63^{\circ}$. Here and below colors indicate the total number of photon counts registered by a pixel. Numbers along the axes represent pixel numbers.
@@
Now let us consider reconstruction of arrival directions of UHECRs. Figure 5 presents the distribution of errors in the angular separation of true and predicted arrival directions for a test sample consisting of 200 events. The mean angular separation equals $4.1^{\circ}$ with the median equal to $3.3^{\circ}$. The coefficient of determination equals 0.948.222In some tests, median of the angular separation was below $1.5^{\circ}$ but we are presenting a more typical result.
@@
It is interesting to notice that, according to our preliminary tests, the above results can be improved if the input data is arranged in heaps of “screenshots” of the focal surface instead of just integral tracks. One can say that in this case the CNN receives multi-channel images instead of “black-and-white” ones. This representation of data puts higher demands on computer resources (mainly the size of memory) but allows one to demonstrate kinematics of signals. In particular, the CNN trained to reconstruct ADs on data with 6 channels reduced errors of angular reconstruction for the events shown in the left and right panels of Fig. 7 down to $1.9^{\circ}$ and $10.9^{\circ}$ respectively. However, some outliers remained for this data representation, too. We plan to study this in more detail in the future.
@@ 4.2 EUSO-TA
Another problem is specific to a small telescope with a narrow FoV, like EUSO-TA. The point is that it is able to observe only a small part of a shower, which results in incomplete information about the shower luminosity and, thus, energy; see an in-depth discussion of this problem in [10]. Thus it didn’t come as a surprise when our early attempts to reconstruct energy basing on pure integral tracks failed. However, arranging input data in heaps of “screenshots” of the focal surface greatly improved the situation. In what follows, we present results obtained for data in which every event was presented by 12 “screenshots” of the FS made in consecutive moments of time. [^3] Data with the time resolution of 1 $\mu$s (or another one) might need another number of time steps to be used.
@@
Figure 10: True (circles) and predicted values of energy for a test sample for EUSO-TA. Red diamonds indicate predictions that deviate from true labels by less than 10%.
@@ 5 Track recognition
Before trying to recognize a track, one has to decide how to label pixels for training the model. In our approach, every data sample (an integral track for EUSO-SPB2 or a 12-layer data chunk for EUSO-TA), was scaled to $[0,1]$ before processing. This allowed us to treat dim and bright pixels on the same ground. Then we considered how we should label pixels on data samples without background illumination to reach the best recognition of hit pixels in noisy data. We tested several thresholds $\alpha=0.1,\dots,0.5$ assigning all pixels above the threshold a value 1 and 0 to all the rest, training the model and then evaluating the performance of track recognition. Figure 17 presents results of one of such tests.
@@
To perform the comparison shown in Fig. 17, we used a sample of events generated for 50 EeV proton primaries for EUSO-SPB2 and tested it on a sample consisting 1000 events. Training of the model was repeated 10 times for each cut with different initial weights. One can see that all three metrics demonstrate a local maximum in the range of thresholds $\alpha\approx 0.25,\dots,0.35$ with slightly lower cuts when training on samples of lower energies. Thus, in what follows we mark hit pixels as belonging to a track if the value of the scaled signal is above 0.25.
Let us illustrate how the implemented procedure works using a “snapshot” of the focal surface of EUSO-TA. The top left panel in Fig. 18 shows a signal recorded in one moment of time. The top right panel shows how the track was marked using the pure signal. The bottom row presents the result of applying the model: the left panel shows predicted probabilities of pixels to belong to the track, and the right one shows the recognized track.
Figure 18: An example of track recognition of EUSO-TA. The top row: the original signal (left) and the labels assigned using the signal without background illumination (right). The bottom row: predicted probabilities for the pixels to belong to the track (left) and the assigned labels (right). See the text for details.
@@ 6 Discussion
We have presented two proof-of concept neural networks that can be used to recognize tracks of extensive air showers registered by fluorescence telescopes and to reconstruct energy and arrival directions of primary UHECRs. It is clear that final results of reconstruction will be worse than those presented above because we have considered a simplified model. In particular, our preliminary tests performed for the EUSO-SPB2 data demonstrated that mean absolute percentage error of energy reconstruction increases by approximately 4% if only hit pixels recognized with mean IoU $\geq 0.89$ are taken into consideration. Still, the accuracy of energy reconstruction remains comparable with estimations for a much more sophisticated JEM-EUSO telescope obtained by conventional methods [14]. In general, it is clear from the above results that special care should be taken for events that have just a few hit pixels, those with tracks touching only the very edge of the focal surface or located along the gaps between PDMs (in the case of EUSO-SPB2). However, this is equally true for conventional methods.
@@ Acknowledgements.
The development of neural networks for EUSO-SPB2 was partially supported by the RNF grant 22-22-0367. Simulations and the development of neural networks for EUSO-TA is supported by the RNF grant 22-62-00010.

### v2

@@ Abstract
Fluorescence telescopes are important instruments widely used in modern experiments for registering ultraviolet radiation from extensive air showers (EASs) generated by cosmic rays of ultra-high energies. We present proof-of-concept convolutional neural networks aimed at reconstruction of energy and arrival directions of primary particles using model data for two telescopes developed by the international JEM-EUSO collaboration. We also demonstrate how a simple convolutional encoder-decoder can be used for EAS track recognition. The approach is generic and can be adopted for other fluorescence telescopes.
@@ 1 Introduction
Registering fluorescence radiation of extensive air showers (EASs) with dedicated telescopes in nocturnal atmosphere of Earth is an established technique of studying ultra-high energy cosmic rays ($\gtrsim 10^{18}\leavevmode\nobreak\ \text{eV}=1\leavevmode\nobreak\ \text{EeV}$[1]. Fluorescence telescopes (FTs) play an important role in both leading experiments in this field of astrophysics, the Pierre Auger Observatory [2] and the Telescope Array [3]. They were also suggested to be used as orbital telescopes due to the opportunity to increase the exposure of an experiment dramatically [4]. The international JEM-EUSO collaboration is implementing a step-by-step program to conduct a full-scale orbital experiment aimed at solving the long-standing puzzle of the nature and origin of the highest energy cosmic rays [5, 6, 7]. Two of the instruments already built are the small ground-based EUSO-TA telescope [8, 9, 10], which operates at the site of the Telescope Array experiment, and the EUSO-SPB2 stratospheric experiment, which included both fluorescence and Cherenkov telescopes [11, 12].
Reconstruction of parameters of initial UHECRs based on the signal registered by a monocular FT is not a trivial task [13, 14, 15]. Here we suggest two simple artificial neural networks that can be used to recognize tracks registered by a single FT and to reconstruct energy and arrival directions (ADs) of ultra-high energy primaries. We employ data simulated for EUSO-TA and the fluorescence telescope of EUSO-SPB2 to train and test these networks. In what follows, we present preliminary results of this technique and outline possible ways of its improvement. We believe that the suggested approach is generic and can be adopted for other fluorescence telescopes operating nowadays as well as the future ones. It can also be modified to use benefits provided by stereoscopic observations.
@@ 2 Fluorescence telescopes EUSO-SPB2 and EUSO-TA
EUSO-SPB2 (the Extreme Universe Space Observatory on a Super Pressure Balloon 2) was a stratospheric pathfinder mission for a future orbital experiment like POEMMA [16]. The scientific equipment included two telescopes with an identical modified Schmidt design of optics with a 1 m diameter entrance pupil occupied by an aspheric corrector plate, a curved focal surface (FS), and a spherical primary mirror. The Cherenkov telescope was designed to observe Cherenkov emission of cosmic ray EASs with energies above 1 PeV. It could be tilted up and down between $3.5^{\circ}$ and $-13^{\circ}$ w.r.t. the horizon pointing towards the limb [17, 18]. The fluorescence telescope was intended to register the fluorescence emission of extensive air showers produced by UHECRs with the energy above 2 EeV. The telescope was pointed in the nadir direction and had a total field of view (FoV) of $12^{\circ}\times 36^{\circ}$. The FT camera consisted of three photo detector modules (PDMs), each built of $48\times 48$ pixels in a rectangular grid, thus giving 6912 pixels in total. Every PDM was covered with a BG3 filter to limit its sensitivity to the near-UV range (290–430 nm). The time resolution was equal to 1 $\mu$[19, 20].
@@
EUSO-TA can operate at different elevation angles. In what follows, the elevation angle of $10^{\circ}$ was used. Studies for different other elevation angles are to be performed in the future.
@@ 3 Simulated data sets
We used CONEX [21] to simulate EAS from UHECRs in the energy range from a few EeV up to 100 EeV. The response of both instruments was simulated with the $\overline{\rm Off}$$\underline{\rm line}$ framework [22]. For EUSO-TA, we demanded that EAS cores were within the projection of the telescope FoV on the ground. The distance between the telescope and shower cores varied in the range 2–40 km depending on the energy of primary protons. For both instruments, we used events that flagged software triggers of the respective telescope. Out of these events, we excluded those that contained just a few hit pixels in one of the corners of the field of view of a telescope and less than two time frames with hit pixels. [^1] The right panel of Fig. 16 provides an example of a signal that was considered for reconstruction despite of the small number of hit pixels. No other quality cuts on the signals were implied. In particular, we did not select events with the shower maximum being in the FoV of a telescope, which can considerably improve the quality of reconstruction.
@@ 4.1 EUSO-SPB2
Figure 1: Examples of integral tracks of simulated EASs produced by UHECRs as seen by EUSO-SPB2. A track shown in the left panel was produced by a shower initiated by a 96 EeV proton arriving at the zenith angle $\theta=19^{\circ}$. The right panel shows a track from a 26 EeV proton arriving at $\theta=63^{\circ}$. Here and below colors indicate photon counts per pixel. Numbers along the axes represent pixel numbers.
@@
Now let us consider reconstruction of arrival directions of UHECRs. Figure 5 presents the distribution of errors in the angular separation of true and predicted arrival directions for a test sample consisting of 200 events. The mean angular separation equals $4.1^{\circ}$ with the median equal to $3.3^{\circ}$. The coefficient of determination equals 0.948.333In some tests, median of the angular separation was below $1.5^{\circ}$ but we are presenting a more typical result.
@@
It is interesting to notice that, according to our preliminary tests, the above results can be improved if the input data is arranged in stacks of “screenshots” of the focal surface instead of just integral tracks. One can say that in this case the CNN receives multi-channel images instead of “black-and-white” ones. This representation of data puts higher demands on computer resources (mainly the size of memory) but allows one to demonstrate kinematics of signals. In particular, the CNN trained to reconstruct ADs on data with 6 channels reduced errors of angular reconstruction for the events shown in the left and right panels of Fig. 7 down to $1.9^{\circ}$ and $10.9^{\circ}$ respectively. However, some outliers remained for this data representation, too. We plan to study this in more detail in the future.
@@ 4.2 EUSO-TA
Another problem is specific to a small telescope with a narrow FoV, like EUSO-TA. The point is that it is able to observe only a small part of a shower, which results in incomplete information about the shower luminosity and, thus, energy; see an in-depth discussion of this problem in [10]. Thus it didn’t come as a surprise when our early attempts to reconstruct energy based on pure integral tracks failed. However, arranging input data in heaps of “screenshots” of the focal surface greatly improved the situation. In what follows, we present results obtained for data in which every event was presented by 12 “screenshots” of the FS made in consecutive moments of time. [^4] Data with the time resolution of 1 $\mu$s (or another one) might need another number of time steps to be used.
@@
Figure 10: True (circles) and predicted values of energy for a test sample for EUSO-TA. Diamonds indicate predictions that deviate from true labels by less than 10%.
@@ 5 Track recognition
Before trying to recognize a track, one has to decide how to label pixels for training the model. In our approach, every data sample (an integral track for EUSO-SPB2 or a 12-layer data chunk for EUSO-TA), was scaled to $[0,1]$ before processing. Then we considered how we should label pixels on data samples without background illumination to reach the best recognition of hit pixels in noisy data. We tested several thresholds $\alpha=0.1,\dots,0.5$ assigning all pixels above the threshold a value 1 and 0 to all the rest, training the model and then evaluating the performance of track recognition.
We considered samples simulated for the whole range of energies as well as samples with a fixed energy of the primary particle. Figure 17 presents results obtained for tracks generated by 50 EeV protons. The whole sample consisted of 14.6 thousand events. Training of the model was repeated 10 times for each cut with different initial weights. For each run, the initial sample was split in a random fashion into 1000 test events and 13.6 thousand events used for training. One can see that all three metrics demonstrate a local maximum in the range of thresholds $\alpha\approx 0.25,\dots,0.35$. Similar results were obtained in other tests. Thus, in what follows we mark hit pixels as belonging to a track if the value of the scaled signal is above 0.25.
@@
Let us illustrate how the implemented procedure works using a “snapshot” of the focal surface of EUSO-TA. The top left panel in Fig. 18 shows a signal recorded in one moment of time. The top right panel shows how the track was marked using the pure signal. The bottom row presents the result of applying the model: the left panel shows predicted probabilities of pixels to belong to the track, and the right one shows the recognized track with the probability threshold equal to 0.5.
Figure 18: An example of track recognition for EUSO-TA. The top row: the original signal (left) and the labels assigned using the signal without background illumination (right). The bottom row: probabilities predicted by the neural network for the pixels to belong to the track (left) and the assigned labels (right). See the text for details.
@@ 6 Discussion
We have presented two proof-of-concept neural networks that can be used to recognize tracks of extensive air showers registered by fluorescence telescopes and to reconstruct energy and arrival directions of primary UHECRs. It is clear that final results of reconstruction will be worse than those presented above because we have considered a simplified model. In particular, our preliminary tests performed for the EUSO-SPB2 data demonstrated that mean absolute percentage error of energy reconstruction increases by approximately 4% if only hit pixels recognized with mean IoU $\geq 0.89$ are taken into consideration. Still, the accuracy of energy reconstruction remains comparable with estimations for a much more sophisticated JEM-EUSO telescope obtained by conventional methods [14]. In general, it is clear from the above results that special care should be taken for events that have just a few hit pixels, those with tracks touching only the very edge of the focal surface or located along the gaps between PDMs (in the case of EUSO-SPB2). However, this is equally true for conventional methods.
@@ Acknowledgements.
The development of neural networks for EUSO-SPB2 was partially supported by the Russian Science Foundation grant 22-22-0367. Simulations and the development of neural networks for EUSO-TA is supported by the RSF grant 22-62-00010.
CONFLICT OF INTEREST
The authors declare that they have no conflicts of interest.
Publisher’s Note
Pleiades Publishing remains neutral with regard to jurisdictional claims in published maps and institutional affiliations.